Disposal of Fixed Asset

Perfect Auto Rentals sold one of its cars on January 1, 2019. Perfect had acquired the car on January 1, 2017, for $13,500. At acquisition Perfect assumed that the car would have an estimated life of 3 years and a residual value of $3,000. Assume that Perfect has recorded straight-line depreciation expense for 2017 and 2018.

Required:

Prepare the journal entry to record the sale of the car assuming the car sold for (a) $6,500 cash, (b) $4,000 cash, and (c) $6,600 cash. The company recorded the car as equipment. If no entry is required, leave the answer boxes blank.

a.
Record sale of car
b.
Record sale of car
c.
Record sale of car
Answer #1

Car was purchased on 1st January 2017 for $ 13,500

Useful life = 3 years

residual value = $ 3,000

Therefore depreciation for each year = (13,500-3,000)/3 = $ 3,500

Therefore Depreciation from 1st Jan 2017 to 1st Jan 2019 = $ 3,500 * 2 = 7,000

Therefore WDV as on 1st January 2019 = (13,500-7,000) = $6,500

Entry for sale of car

a.) Journal entry when car is sold for $ 6,500

Dr Cash $6,500

Dr Accumulated Depreciation $7,000

Cr Equipment A/c $ 13,500

b.) Journal entry when car is sold for $ 4,000

Dr Cash $4,000

Dr Accumulated Depreciation $7,000

Dr Loss on sale of equipment $2,500

Cr Equipment A/c $ 13,500

c.) Journal entry when car is sold for $ 6,600

Dr Cash $6,600

Dr Accumulated Depreciation $7,000

Cr Equipment A/c    $ 13,500

Cr Profit on sale of equipment $100